OpenStudy (anonymous):

At maximum speed, an airplane travels 1680 miles against the wind in 5 hours. Flying with the wind, the plane can travel the same distance in 4 hours. Let x be the maximum speed of the plane and y be the speed of the wind. What is the speed of the plane with no wind?

OpenStudy (anonymous):

let , m is the speed of plane with no wind 1680/5=336 1680/4=420 x=420=m+y m-y=336 hence, 2*m= 756 hence,m=378 miles/ hour

OpenStudy (anonymous):

so with no wind its 378?

OpenStudy (anonymous):

yes i think so..

OpenStudy (anonymous):

yea i got it right thank you

OpenStudy (anonymous):

you are most welcome..

OpenStudy (anonymous):

can u help me with another question that was like the last one i posted up

OpenStudy (anonymous):

can u paste the link?..sry i cannt find..

OpenStudy (anonymous):

i just posted it up

OpenStudy (anonymous):

ceaser i cannt find the qwestion you are talking about...